OpenStudy (anonymous):

Solve: (x-1)/(x+3)>4

OpenStudy (anonymous):

x-1=4(x+3), solve for x

OpenStudy (anonymous):

x-1>4(x+3) ***

OpenStudy (anonymous):

oops beware!

Directrix (directrix):

-13/3 < x < -3

OpenStudy (anonymous):

it might be easier to see when written as (x-1)/(x+3) = 1 + (-4)/(x+3) > 4 so that values of x less than -3 but close enough will work

OpenStudy (anonymous):

(x-1)/(x+3) - 4 > 0 (x-1)/(x+3) - 4(x+3)/(x+3) > 0 (x-1-4x-12) / (x+3) > 0 (-3x -13) / (x+3) > 0 x cant be -3 because it will make the denominator zero. then plug any value of x : * between -3 and -13/3 (the value of the left side will be positiive) *larger than -3 (the value of the left side will be negative) *and smaller than -13/3. (the value of the left side will be negative since (-3x -13) / (x+3) > 0 , we need the value of the left side to be positive. therefore the solution is -13/3 < x < -3

OpenStudy (anonymous):

Thank you!